OpenStudy (anonymous):

Please help Use the Substitution Method to solve the following system of equations. x - 4y = 1 2x - 9y = 3 (-1, -3) (-3, 1) (3, 1) (-3, -1)

OpenStudy (anonymous):

A B C or D?

OpenStudy (ledah):

use the method it told you.. as you know. we cannot just give out the answers.

OpenStudy (anonymous):

why not?

OpenStudy (anonymous):

get one equation to be in y=mx+b form

OpenStudy (anonymous):

i pick the first equation

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now put it in mx+b form

OpenStudy (anonymous):

then subsutite y for the second equation

OpenStudy (ledah):

we cannot give out the answers only because we aren't Question and Answer site.

OpenStudy (anonymous):

4y-x=1

OpenStudy (anonymous):

ledah... we cud help him get the answer thou right?

OpenStudy (anonymous):

is there a questions and answers site?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

its the last one

OpenStudy (ledah):

Yes we can. help them get the answer. You just cannot give them the answer..

OpenStudy (ledah):

By just giving out answers.. The student doesn't learn anything

OpenStudy (anonymous):

Part 1: Solve the following system of equations. Please show your work to receive full credit. x - y = 16 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice.